use the given conditions to write an equation in the point-slope form and the slope-intercept form
slope = 8, passing through (4, -1)

Answer by venugopalramana(3286)

use the given conditions to write an equation in the point-slope form and the slope-intercept form
slope = 8 = M, passing through (4, -1)= (X1,Y1)
FORMULA IS
Y-Y1=M(X-X1)
Y-(-1)=8(X-4)
Y+1=8X-32
Y=8X-33......................SLOPE = 8....INTERCEPT = -33